OpenStudy (anonymous):

any1 know how to find the amount of the variable when working with percents using 6th grade math?

OpenStudy (amistre64):

an example might be useful to work with

OpenStudy (anonymous):

364 is what percent of 455

OpenStudy (amistre64):

1 percent is equal to a division by 100 1% of 455 = 455/100 how many 1% are in 364? divide 364 by the value of 1%

OpenStudy (anonymous):

all i need to know is how to set the problem up. my teacher told us that we need to use cross multipling

OpenStudy (anonymous):

btw this is mt h/w that i cant figure out

OpenStudy (amistre64):

divide 364 by some unknown value: x and equate it to that 455/100 \[\frac{364}{x}=\frac{455}{100}\] that works out to the same thing .... just solve for x

OpenStudy (anonymous):

exactly what i need to know

OpenStudy (amistre64):

there are actually a number of different ways to set up that same problem ... \[\frac{364}{x}=\frac{455}{100}\] \[\frac{364}{455}=\frac{x}{100}\] \[\frac{100}{455}=\frac{x}{364}\] \[\frac{100}{x}=\frac{455}{364}\] all of these would produce the same result

OpenStudy (amistre64):

the second one is prolly the "correct" setup if you want to go by how textbooks do it

OpenStudy (anonymous):

we dont use textbooks in my class

OpenStudy (anonymous):

100 times ? is 100? and 364 times 100 is 36400 then i have to divide

OpenStudy (anonymous):

? is the variable

OpenStudy (anonymous):

so it would be